### v1

# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

Abstract
Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community, unlike previous virtual worlds. Its users can create and publish wearables, virtual garments to customize avatars, which can be then sold or given away via the blockchain. Decentral Games (DG), a single project owning two prominent in-world casinos, has by far created the most wearables, with these being necessary to earn cryptocurrency in their flagship game ICE Poker. We thus present a comprehensive study that investigates how DG and ICE Poker influence the overall dynamics of Decentraland wearables and in-world visits. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that the influence of DG and Ice Poker is not only significant, but also substantial for transfers and sales monetary value of wearables, and very large for daily unique visitors and time spent in the virtual world. Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, online poker appears to be the main driver of the analyzed dynamics. Our work thus contributes to the current understanding of user behavior in social virtual worlds and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual spaces.
keywords:
social virtual worlds, Metaverse, online gambling, NFT, blockchain
1 Introduction
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision that is in its early stages Cheng et al. (2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022). In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al. (2022). In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user. Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019). However, it has been argued Zhou et al. (2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are detrimental to users due to the imbalance in terms of rights between platforms and users. For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the platform ceases to function.
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020). Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a more scalable and compatible blockchain. Hence, there is a trove of public data regarding the dynamics of this social virtual world.
Based on these data, Trujillo and Bacciu 2023 were among the first to conduct a quantitative study of Decentraland’s wearables. In their analysis, it emerged that a single account —related to the project Decentral Games (DG)— was by far the one with most published wearables. DG manages a few in-world virtual casinos, and possession of DG wearables is a requirement to earn money (in the form of cryptocurrency) by playing ICE Poker, the main project’s attraction (depicted in Figure 1). In addition, owners can delegate their DG wearables to other users, with whom earnings are split based on the wearable’s level, which can be also upgraded for a fee. Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing to users of Decentraland due to their earning potential.
Figure 1: Screenshot of an ICE Poker game session inside a Decentral Games casino on Decentraland. Note the avatar wearing the sleeveless shirt with the name of the game, which allows the player to earn cryptocurrency, as well as the croupier being based on the Internet meme Pepe the Frog.
Therefore, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland, a virtual world that touts its social and creative aspects of all kinds, but that seems to be driven by gambling. In particular, we seek to answer the following research questions (RQs):
RQ1:
Are transfers of Decentraland wearables disproportionately related to those created by Decentral Games?
RQ2:
Do Decentraland users visit significantly more ICE Poker casinos compared to other in-world locations?
To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of transferring ownership of Decetranland’s wearables in its different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers. To answer RQ2, we quantify the visits —both in time and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land. We also consider and describe the time series characteristics of these visits on a daily basis.
1.1 Contributions and significance
Our main contribution is the first quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world. Thanks to its trailblazing development based on blockchain and tokenomics, Decentraland gives a glimpse of how new kinds of immersive social media might work on the so-called Metaverse. Therefore, studying the use (and potential abuse) of the platform’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds.
2 Background
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and DG referred to in the related work, analyses, and discussion thereof.
2.1 Decentraland
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020). On Ethereum and Polygon, these exchanges are made from one address to another address via smart contracts, which set the automatic rules of the transactions. These addresses identify either an account or a smart contract that can participate on the blockchain.
Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables. The proof of ownership of these assets is in the form of non-fungible tokens (NFTs) —unique and non-interchangeable, unlike the fungible cryptocurrencies. Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization (DAO) of Decentraland, a form of blockchain-based governance without a central authority Goldberg & Schär (2023).
The world (see Figure 2) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid, whose coordinates range in both axes from -150 to 150. Land is divided into two main types: public and privately owned. Public land is in turn divided into roads (10.4% of the world map) and plazas (4%), and cannot be sold. Before the public release of the project, 37.4% of the world map was allocated for districts, i.e., groups of adjacent parcels of privately owned land that share similar user interests. District parcels were assigned via an auctioning process to different groups, which respectively manage the land within their boundaries. Among the biggest districts we find, for example, the gambling-focused Vegas City and the shopping-oriented Fashion Street.
Figure 2: Decentraland world map, where the biggest districts are indicated in violet and roads in gray.Vegas City district is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games. They cover only a minuscule surface of the map, accounting for less than 0.1%.
Decentraland wearables are virtual garments, accessories or full body costumes (called skins) that can be used to clothe one’s avatar or completely change its appearance. Wearables are grouped into collections of one or more items. Each item represents a distinct design from which one or more NFTs can be minted, i.e, a new unique token ID representing the asset is created and assigned. Mints are registered on the blockchain as transfers from a special null address. In addition, a given item has a rarity limit (from a predefined list) that sets its maximum supply of tokens —e.g., a unique item can be minted only once, while a common item 100,000 times. To publish a wearable collection, creators pay a fee and submit their designs to the approval committee appointed by the DAO. Once approved, creators can sell their wearables via Decentraland’s Marketplace (for a fee that goes to the DAO), or they can mint and transfer the respective NFTs to a blockchain address (their own or not), either as a gift or to be sold on a different marketplace.
Originally, in the first version of wearables, these were created via a contest with approved designers and minted on Ethereum. However, this proved to be too limiting and prohibitive due to the then high transaction fees on that blockchain. Therefore, the wearables system was changed to the second and current version described above, now open to all and based on the less costly and more scalable Polygon blockchain. For the launch of both versions, a fashion event was held to showcase this creative endeavour and promote the newly published wearables; these events would be the precursors of more elaborate fashion initiatives. Arguably, the most important subsequent initiative was the first edition of the Metaverse Fashion Week in March 2022, a series of in-world social events that involved renowned fashion brands such as Tommy Hilfiger and Dolce & Gabbana.
2.2 Decentral Games
The project DG started in 2019, a year prior to Decentraland’s public release, and has played a part in many initiatives in the virtual world ever since. In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was one of the few creators involved in the short-lived first version of Decentraland wearables. DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance.
In October 2021, DG released ICE Poker, touted to be a Metaverse play-to-earn game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level. These were made available for minting in three batches to users that had a certain amount of DG and ICE.
Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas District. The first one and more popular is called The Stronghold; the second one, Diamond Hands City, is exclusive to accounts that possess ICE-themed wearables with the highest rank. Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, now called ICE Poker Arcade. In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or virtual assets such as wearables.
The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure 3; the more expensive the better. Once inside either game mode, players can wear up to five wearables to augment their rewards. For certain ICE-themed wearables, it is also possible to upgrade their ranking level (ranging from 1 to 5) to gain better rewards by paying a smaller fee compared to buying one. Additionally, all DG wearables can be delegated to other users, who can then play and split their earnings with the owner of the wearable, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
Figure 3: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus in ICE Poker; b) DG Suited Tracksuit Bottom, which can be used to play ICE Poker Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus in ICE Poker.
3 Related Work
While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009). World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010). Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al. (2018).
In response to these challenges, a wave of blockchain-based virtual experiences has emerged Cannavo & Lamberti (2020); Huynh-The et al. (2023). Virtual worlds such as Decentraland Ordano et al. (2017), The Sandbox Madrid et al. (2020), Axie Infinity Nguyen et al. (2021), Upland Honigstein et al. (2019), and Illuvium Warwick & Warwick (2023) all attempt to mitigate the aforementioned issues by leveraging the main strengths of blockchain technology (transparency, security, and decentralization) for various purposes. These include verifying the digital ownership of virtual assets, enabling cryptocurrency-based economies, and implementing decentralized governance systems.
In particular, Decentraland constitutes a standout example, being the earliest and most popular virtual world based on blockchain. It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena in these novel asset-oriented social media. However, existing research on Decentraland has primarily focused on landholding assets Dowling (2022); Goldberg et al. (2021); Guidi & Michienzi (2022) or examined the platform’s tokens within the broader context of the NFT ecosystem Nadini et al. (2021). There is a lack of studies regarding the other main kinds of digital assets available in Decentraland, namely usernames and wearables for avatars, with the latter having attracted the attention of the fashion industry Gonzalez (2022).
Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars Bardzell et al. (2010). Despite criticisms that it did not live up to public expectations Rodriguez Sanchez & Garcia-Badell (2022), Second Life maintains a robust economy centered on avatar customization, with many users becoming in-world fashion designers, creating clothing and accessories for avatars Boellstorff (2015).
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022). Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al. (2022), participating in events like the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022). Among the scarce scholarly investigation into Decentraland’s wearables, we find a quantitative analysis in which it emerged that they exhibit peculiar dynamics Trujillo & Bacciu (2023). For instance, wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives. Only a minor portion of them are actually sold on the platform’s marketplace. Furthermore, the price of these wearables is largely influenced by their pre-established rarity, which escalates the cost even further for unique items (i.e., having a set minting limit of one). The study also identified a specific account that stands out in terms of wearable items collections, having published more than 80 at the time of the study. This is noteworthy, given that over half of all wearable creators in Decentraland typically author only a single collection. This creator is associated with Decentral Games, a company whose primary offering is a poker gaming experience within Decentraland virtual casinos.
Online gambling is a topic of research in its own right. This includes studies on the legality of online gambling and cryptocurrencies in different jurisdictions Zborowska et al. (2012); Yeoh (2017), the role of virtual communities in either motivating gamblers or protecting them from excessive behaviour Sirola et al. (2021); Savolainen et al. (2022), the overlap of personality and demographic traits among cryptocurrency traders and problem gamblers Johnson et al. (2023); Delfabbro et al. (2021), how those traits are partly different in professional gamblers Newall & Talberg (2023), and how playing habits in low-stakes gambling might influence the perception of financial risk in general Armstrong et al. (2018).
4 Methods
For our study, we focus on the two years after the launch of Decentraland wearables on Polygon, that is, from June 2021 to May 2023. We excluded the short-lived first version of wearables because these were already deprecated by the launch of ICE Poker. In the following, we describe the overall approach to answer our RQs and then we detail how we performed the respective data collection.
4.1 Overall approach
For RQ1 (wearable transfers) we first retrieved all wearables created during the period of study, marking those created by DG based on metadata from Decentraland. We then retrieved from the Polygon blockchain all the transfers related to all wearables, as well as their sales made in the most prominent marketplaces. We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers as well as sales volume (both in number and USD). We also analyzed the share of transfers related to upgrades, which we retrieved from the respective log events on Polygon.
For RQ2 (in-world visits to DG poker casinos), we first collected in-world tracking data of users, available only from 12 July 2022. Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which DG poker casinos are located, compared to the rest of Decentraland’s map.
Our descriptive and statistical analyses mostly concern proportions, i.e., the share of wearables or land associated with DG compared with the rest of Decentraland. For RQ1, this is either for the whole two-year period or on a monthly basis therein. For RQ2, this is either for a ten-month period or on a daily basis, as we have a narrower time frame but a finer time resolution for data of in-world visits. To test for significant differences in proportions we used two-sided tests with significance level $\alpha=0.05$. When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure the effect size. For daily in-world visits, we also used iterative STL (i.e., seasonal and trend decomposition using LOESS) to analyze the time series, including trend and seasonal strengths (from 0 to 1).
4.2 Data collection
For wearable metadata (e.g., collection, category, creator), we used several of the openly available Decentraland’s query endpoints on TheGraph, a service for querying blockchain networks, and whose definitions are open source. [^1] https://github.com/decentraland For transfers (including mints), we collected the respective transactions from Polygon via the blockchain web service Alchemy. For wearable sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
We also retrieved from Yahoo! Finance the price of the few different cryptocurrencies used for the collected wearable sales during the two-year period of study. To normalize the price of sales, we first converted it to USD, then we adjusted it for inflation on a monthly basis as of December 2023, based on the Consumer Price Index (CPI) by the Bureau of Labor Statistics of the United States. Price changes of these cryptocurrencies also serve to gauge the interest on the studied platforms. Indeed, as can be seen in Figure 4, the ICE token was released around Meta’s rebranding in November 2021, after the latter provoked a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days). This was a considerable surge, even taking into account the already upward trend of MATIC in the wider Polygon ecosystem. However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter OECD (2022), which affected a large share of blockchain projects and lasted until mid 2023.
Figure 4: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract which manages the respective on-chain transactions, called “iceRegistrant”. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant from Decentraland’s DAO to finance a system to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
5 Analyses and Results
For the two-year period of study, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
5.1 RQ1: Overall wearable transfers
[…]
In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but interestingly, the variations are different for transfers and sales. On one hand, and as illustrated in Figure 5, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March. In those two months the share of DG wearables was only 3.3% and 4.5%, respectively. However, the transfers monthly share of DG wearables increased noticeably from September 2022 onward, reaching its peak in May 2023 with 85.3%, while the rest of wearable transfers decreased in number. This DG share increase is mostly related to a few wearable items with a common rarity limit and which give access only to ICE Poker Arcade, whose app was released precisely at the end of September. On the other hand, for the monthly sales volume in USD (illustrated in Figure 6), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting. Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure 4.
Figure 5: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and respective share associated with Decentral Games (DG). The upward trend in the bottom panel is related to wearables that grant access to the ICE Poker Arcade app, released in September 2022.
Figure 6: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG). The bottom panel’s spike in November 2021 is related to the release of ICE-themed DG wearables and the high valuation of ICE in the wake of Meta’s rebranding.
For the two-year period there were 58,791 mints of wearables with an ICE ranking level, of which 42,390 (72.1%) were via upgrades. Each upgrade implies three wearable NFT transfers on Polygon: 1) transfer of the original NFT from the owner to the upgrading smart contract; 2) burning of this NFT by transferring it to a special dead address, effectively removing it from circulation; and 3) minting a new NFT for the owner from a higher-ranking wearable item within the same collection. Hence, there were 127,170 transfers related to upgrades, representing 14.2% of DG wearable transfers and 2.1% of all transfers. Interestingly, only 10% of the mints of ICE-ranking wearables with a level higher than 1 was not related to an upgrade. In other words, 90% of owners preferred to burn a lower-ranking token and mint a higher-ranking token via an upgrade, rather than buy (or wait to win) one without sacrificing the former.
5.2 RQ2: In-world visits of poker casinos
[…]
When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figurge 7) and number of distinct visitor accounts (see Figure 8). For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors, with a trend strength (based on STL decomposition) of 0.98 and 0.99, respectively. In the rest of the world map, the trend for time spent is slightly downward and with a strength of 0.83 for time spent, while the trend is visibly downward with a strength 0.91 for visitors. In other words, over time there were less distinct accounts visiting Decentraland, but these collectively spent more time on it. However, upon further inspection this increase in time spent per visitor is mostly due to a few accounts with a much higher number of hours than average. Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due the the aforementioned outlier users in the rest of the world map. Nevertheless, for visitors, the share at the beginning of the ten-month period was of circa 32%, and at the end it increased to circa 41%. Curiously, the number of visitors of ICE Poker casinos had a modest weekly seasonality strength of 0.66, with the number of visitors slightly decreasing on weekends; for the other time series the seasonality strength was less than 0.5.
Figure 7: Stacked area chart of daily time spent (in hours) visiting Decentraland. ICE Poker casinos had a remarkable share, with a daily median of 20%. The highest overall spike was reached during the Metaverse Music Festival (MVMF) in mid-November.
Figure 8: Stacked area chart of daily distinct accounts that visited Decentraland. ICE Poker casinos had a striking share, with a daily median of 33%. The Metaverse Music Festival (MVMF) in mid-November had the highest spike overall, while the Metaverse Fashion Week (MVFW) at the end of March had the only spike during the last four months.
These trend and seasonality strength differences and shares by land category are in part explained by their dissimilar spikes and surges. On one hand, visits to the ICE Poker casinos had no spikes nor dips. There was, however, a short-lived surge (for both time and accounts) that started at the end of August 2022, when DG gifted special wearables to circa 3,000 of the most active players as a reward and in preparation for its new Tournament game mode. On the other hand, in the rest of the world map there were several visit spikes, surges, and dips. In particular, the Decentraland Metaverse Music Festival, held on 10–13 November 2022, was the event that provoked the most visits (both in time and accounts). The festival was free to attend, and its line up involved famous artists such as Björk, Ozzy Osbourne, Dillon Francis, and Soulja Boy. Incidentally, the Metaverse Fashion Week of 2023 also provoked a spike in visitors at the end of March, but not in time spent and only during a period of four months in which the number of distinct visitors had already remarkably declined.
6 Discussion
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we first discuss our main findings regarding the influence of DG and ICE Poker on Decentraland, we comment on the in-world economic mechanisms that allowed such influence to exist, we reflect upon the ambiguous legal standing of ICE Poker, and, finally, we describe the main limitations of our study.
6.1 Main findings
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Still, this influence has varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
The Metaverse Fashion Week is arguably the most important initiative around Decentraland wearables. Both editions of 2022 and 2023 attracted well-established fashion brands, were featured in mainstream media, and piqued the interest of the general public. The initiative, however, did not meet expectations. Based on a qualitative study on the 2023 edition Miikkulainen-Gilbert (2023), most user complaints were related to the on-boarding complexity of setting up an Ethereum account, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction for many events, the technical limitations of wearable 3D modeling, and the need of paid approval even for wearables designed for oneself. In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure 8), but there was no visible effect on the time spent on the platform (see Figure 7), and the number of visitors continued its downward trend afterward. The fact that in 2024 there was no Metaverse Fashion Week confirms the failure of the initiative to keep up the momentum going. On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures 7 and 8), but nevertheless visits did not visibly increase in the following weeks. Unlike in-world fashion events, music concerts have been present for many years in social virtual worlds and MMO games, and during the Covid-19 pandemic they were seen as an attractive alternative for the struggling live music industry Çeli̇k (2023). It remains to be seen if this Decentraland initiative will continue to gather interest in future editions.
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation. Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023). In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables, however, is still necessary. This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023). Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6.
Nevertheless, the interest on the Metaverse has now greatly subsided.
Over time, these platforms have tried to adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations.
Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024). DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
6.2 On the legality of ICE Poker
After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions Zborowska et al. (2012). Similarly, for many years and due to their novelty, blockchain cryptocurrencies remained in a regulatory limbo in most jurisdictions Yeoh (2017). However, after many instances of scams and fraud —some very large and highly mediatized, as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies Trautman et al. (2022). ICE Poker is a combination of both gambling and cryptocurrencies, hence, this begs the simple question on whether it is a legal operation or not. The answer is, however, murky.
Firstly, on the project’s front page it is stated that DG is operated by the company Kidderminster Cheltenham, whose official address is in Costa Rica. This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  Kerr (2018), its tax haven tendencies Baron (2019), and the fact that its operating licenses are often used to offer an online gambling service in foreign jurisdictions in which the licensing process is more rigid or non-existing Fischer (2021). Moreover, in the official project documentation it is stated that the company behind DG is called Web4 LTD, which is established in the British Virgin Islands, the place of jurisdiction regarding the terms and conditions of the project, and a British overseas territory known for being a tax haven and having links to dubious online gambling operations Macegoniuk (2021).
Curiously, in the project’s website there is also a legal opinion regarding ICE Poker by a consultant from the Singaporean firm Jacque Law Zhirong (2022), in which it is stated that the game “would not constitute ‘gambling’, ‘betting’, ‘lottery’ under the RGA [Singaporean Remote Gambling Act of 2014]” because “gambling laws are not intended to cover games whereby there is no staking or wagering of money or money’s worth”. Indeed, possession of a DG wearable is necessary to earn ICE, but the wearable has independent value outside of ICE Poker and its use is not exclusive to DG. In addition, users play with chips, which according to the opinion Zhirong (2022), “are non-transferable tokens that hold no value or functionality other than to keep score when players play ICE Poker”. Based on the number of chips and wearables they possess, players periodically receive ICE, which can be converted to fiat money, but on an external exchange such as Binance. This convoluted mechanism is similar to the one used offline by pachinko parlours in Japan to circumvent gambling laws Brooks et al. (2008): pachinko players win small steel balls, which can be exchanged for a receipt with a nominal value, which in turn is used to claim a low-valued physical item (e.g., a specific pack of washing powder), and finally, the player can exchange this item for a lesser nominal value in cash outside the parlour on a nearby “independent office”.
Of course, the above legal opinion applies only to Singapore. In the case of other jurisdictions DG is more cautious, claiming that the onus is on players. In fact, to participate in its virtual casinos, players must guarantee they are not residents of one of the non-supported regions listed on the project’s website. At the moment of writing (early 2024), there were 36 non-supported regions, including high-income countries such as the United States, the United Kingdom, France, Germany, and Australia. ICE Poker Arcade might nonetheless be supported in these regions, given that users play to win physical or digital prizes, not money or cryptocurrency. To address the issue in a more solid manner, recently DG acquired a 90% stake in Virtue Gaming Ltd, the entity that holds a Malta license to enable real-money gambling in Virtue Poker, the first licensed cryptocurrency-based online poker. DG intends to expand this coveted license to its virtual casinos and other games, which would remove doubts regarding their current legal status in many jurisdictions.
6.3 Limitations
In spite of our careful approach, several limiting factors could influence the interpretation and applicability of our results. In particular, despite having public access to all of the wearable transactions on Polygon, we had to make use of sample data for information not readily available in order to better comprehend both transfers and visits.
One such limitation concerns sales data, as we only collected data from two marketplaces. Tracking every single transfer that is also a sale is burdensome because each marketplace has its own smart contracts and sales protocol. Moreover, these might even change over time. For instance, we noticed a sales drop that lasted several days on OpenSea in May 2022, which upon further inspection resulted from a migration to their current protocol called SeaPort. Nonetheless, the marketplaces we selected should cover the vast majority of wearable sales, as Decentraland’s Marketplace is the official and default venue, and OpenSea is the largest NFT marketplace and was the initial default venue to sell wearables by DG, which now has its own marketplace.
Another potential limitation regards in-world tracking of users. We did not use the term active users in our analysis because we only considered the presence of a given account address in the log events, i.e., we did not filter out potential bots or users that could be deemed inactive. In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these Varvello & Voelker (2010). In addition, a user might be authentic (i.e, not a bot) but be away from keyboard while their avatar is still present in the virtual world. Determining an avatar who represents an “authentic active user”, however, depends on an arbitrary set of characteristics and mechanisms. Indeed, calculating the number of daily active users on Decentraland has sparked much public debate, as different definitions and mechanisms might result in vastly different numbers Thompson et al. (2022). Our approach, on the other hand, is relatively simplistic, but at the same time it better responds RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
7 Conclusions
The concept of the Metaverse has in recent years sparked the collective imagination on novel and creative ways in which we could interact within immersive social virtual worlds and represent ourselves via our avatars. Decentraland —whose economy, digital assets, and governance are all based on blockchain technology— offers a glimpse on how a potential Metaverse based on tokenomics might work in the future. Indeed, Decentraland has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds. Moreover, some of its economic and artistic initiatives with worldwide renowned personalities and companies, such as the Metaverse Fashion Week and the Metaverse Music Festival, have reached the mainstream media and gathered much public attention. At present, however, based on our results we can affirm that ICE Poker is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general.